If tan A= 1/3 and tan B = 1/7 then the value of 2A+B=?

Grade:11

1 Answers

Arun
6 years ago
Dear Samira
 

tanA = 1/3, tanB = 1/7

Then tan(2A) = 2tanA/(1-tan^2A) = (2/3)/(1 - 1/9) = 3/4

tan(2A + B) = (tan2A + tanB)/(1 - tan2A.tanB)

tan(2A + b) = (3/4 + 1/7)/(1 - (3/4)(1/7))

tan(2A + B) = (21/28 + 4/28)/(1 - 3/28)

tan(2A + B) = (25/28)/(25/28) = 1

i.e. 2A + B = 45°

 

Regards

Arun (askIITians forum expert)